OpenStudy (anonymous):

how do i find the zeros of f(x) +x (with exponent 2) - 6x +5

OpenStudy (danjs):

factoring, or quadratic formula

OpenStudy (danjs):

f(x) = x^2 -6x + 5

OpenStudy (danjs):

f(x) = (x - 5)(x - 1)

OpenStudy (danjs):

0 = (x-5)(x-1) x-5 = 0 or x-1 = 0

OpenStudy (danjs):

x = 5 or x = 1